OpenStudy (anonymous):

is 4 4/9 divided by 9 2/3 = 2/3 ?

OpenStudy (anonymous):

or 40/87 ?

OpenStudy (anonymous):

4 4/9 = 40/9 9 2/3 = 29/3 (40/9) / (29/3) = (40/9) * (3/29) = 120/261 120/261 = 40/87 (you can factor a 3 out)